OpenStudy (anonymous):

100kg weight object moving with a 8m/s. and it collide with a solid block. how much force will be exert.

OpenStudy (dls):

\[\LARGE F=MA\]

OpenStudy (anonymous):

ya so m=100kg what is A?

OpenStudy (anonymous):

i need answer in newton

OpenStudy (dls):

\[\LARGE V=U+AT\] initial speed=8 final=0 since it comes to rest i suppose do we have time?or distnce?

OpenStudy (anonymous):

time means???? u mean. time needed to stop the object

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Type in the entire question

OpenStudy (dls):

yeah

OpenStudy (anonymous):

its below 1 sec

OpenStudy (anonymous):

0.5 sec

OpenStudy (dls):

v=u+at 0=8+1/2a -8=1/2a a=-16m/s F=ma 100*(-16)=-1600N